On Jan. 1, 2019, ABC Co. issued $52,000 in bonds for $48,000 in cash. The stated interest rate was 4% while the market rate was 8%. The bonds are paid semi-annually on June 30 and December 31 and the bonds mature in 2 years. Journalize the issuance, interest payments, and maturity of the bond.
Answer #1

Required journal entries are:

(a) Issued bonds of $48000 for $52000

Date Description Debit Credit
a Cash 48000
Discount on bonds payable 4000
Bonds payable 52000
(for bond issue at discount)

(b) Semi annual interest payment entry:

Semi annual interest = 4% * $52000 * 6 / 12 = $1040

Assuming straight line method is followed for amortization of bond discount.

Under the straight line method, bond discount is amortized over the life of the bond. Here,

Bond discount = $4000, Life of the bond = 2 * 2 = 4 semi annual years

Straight line amortization of bond discount = $4000 / 4 = $1000

Semi annual cash payment of interest = $1040

Interest expense under straight line method is:

Interest expense = Cash paid for interest + Discount

Interest expense = $1040 + $1000 = $2040

Required journal entry to record the payment of interest:

Date Description Debit Credit
b Interest expense 2040
Cash 1040
Discount on bonds payable 1000
(for interest payment and bond discount amortization)

(c) Journal entry for bond payment at maturity:

Date Description Debit Credit
c Bonds payable 52000
Cash 52000
(for bond payment at maturity)